# Is the System Message Really Important to Jailbreaks in Large Language Models?

### v1

@@ 3.2 Evaluation
    Manual Annotation: Initially, manual annotation (Wei et al., 2023) was relied on, but this method proved to be inefficient.
@@
    Sub-string Matching: This evaluation (Zou et al., 2023; Lapid et al., 2023) is fast, although it suffered from coarse-grained assessments and susceptibility to errors.
@@
    Toxicity: In addition, some researchers (Qi et al., 2023) employed a rater such as Perspective API (Google, 2023b) to evaluate the toxicity of responses, categorizing responses exceeding a predetermined threshold as jailbreak prompts.
@@
    LLMs Evaluation: Another method of evaluation is to utilize the language ability of LLM. Some researchers used LLMs to score responses (Chao et al., 2023), with strict success criteria, raising a greater degree of error rate. Some explored leveraging the few-shot learning capabilities of LLMs chatGLM (Huggingface, 2023a) for prediction (Wang et al., 2020). However, existing studies (Zhao et al., 2021; Lu et al., 2021; Ma et al., 2023) have emphasized the substantial impact of case selection and ordering on performance, which could potentially lead to outcomes close to randomness. Researchers also fine-tune Gopher (Rae et al., 2021) in Bot-Adversarial Dialogue(BAD) (Xu et al., 2021) for evaluation. However, using LLMs for evaluation runs too slow.
  5. 5.
    RobertaYu et al. (2023) evaluated responses using Roberta (Liu et al., 2019) trained on the manually datasets human judgments. They are matched with substrings (Zou et al., 2023), and the jailbreaks are assessed using OpenAI’s various APIs (Markov et al., 2023), including GPT3.5-turbo (OpenAI, 2022) and GPT4 (Bubeck et al., 2023). Their evaluation includes metrics such as True Positive Rate (TPR), False Positive Rate (FPR), and time efficiency. Additionally, they found that the Roberta model demonstrates promising performance in these metrics.
Therefore, we adopt the evaluation methodology proposed by (Yu et al., 2023) to determine whether LLMs have exhibited jailbreak behavior.
4 Jailbreak experiments with different system messages
To determine whether the system message is crucial to jailbreaks in LLMs, in this section we perform the following experiments.
  1. 1.
    First, we apply jailbreak Prompts from models with Short system messages (PS) to Models with Long system messages (ML1) and Models with No system messages (MN1), respectively.
  2. 2.
    Next, we apply jailbreak Prompts from models with Long system messages (PL) to Models with Short system messages (MS1) and MN1, respectively.
  3. 3.
    Finally, we apply jailbreak Prompts from models with No system messages (PN) to MS1 and ML1, respectively.
The results, as depicted in Table 1, illustrate the frequency of prompts leading to successful jailbreaks, with corresponding percentages in parentheses. For instance, under the PS condition, there were 114 (38.0%) and 242 (80.6%) successful jailbreak prompts in the ML1 and MN1 versions respectively. Upon scrutinizing each row of the table, several noteworthy phenomena emerge: (1) Models with different system messages exhibit distinct behaviors in response to the same jailbreak prompts. (2) The PL condition demonstrates strong transferability to MS1 and MN1, while PN exhibits poor transferability to MS1 and ML1. (3) ML1 displays the highest resistance to jailbreak prompts. Even without optimization, ML1 can resist up to 86.4% of PN.
Table 1: jailbreak prompts in model with different system messages
MS1 ML1 MN1
PS –(–%) 114(38.0%) 242(80.6%)
PL 251(83.6%) –(–%) 273(91.0%)
PN 50(16.7%) 41(13.6%) –(–%)
Figure 2: Flow framework diagram of SMEA
To delve deeper into the transferability of jailbreak prompts, we generate synonymous sentences for each system message to obtain LM2, SM2, and NM2, as provided in Appendix A Table 5. These messages are then set as system messages to get models MS2, ML2, and MN2, respectively. Subsequently, we employ PS, PL, and PN to jailbreak MS2, ML2, and MN2, and the results are presented in Table 2.
By juxtaposing the findings from Table 1, we observe that merely substituting system messages with synonymous sentences may render original jailbreak prompts ineffective. Additionally, focusing on the second row of Table 1 and the first column of Table 2, we note that PL exhibits the weakest mobility on ML2 and the strongest mobility on MN1. We term this observation as Security Transferability of System Messages (STSM).
Therefore, we draw the following conclusions:
  1. 1.
    System messages play a crucial role in LLMs jailbreak.
  2. 2.
    Jailbreak prompts exhibit a degree of transferability across different system configurations.
Table 2: Jailbreak for models with synonymous sentences of system messages
@@ 4 Jailbreak experiments with different system messages
203(67.7%) 207(69.0%) 212(70.7%)
5 System Messages Evolutionary Algorithm
Given the profound impact of system messages on jailbreak prompts, we propose System Messages Evolutionary Algorithm (SMEA), devised to procure optimal system messages that mitigate jailbreak risks.
5.1 SMEA framework
The SMEA framework comprises four primary phases: initialization, generation operation, evaluation, and selection.
  1. 1.
    Initialization: In this phase, we leverage the system message commonly utilized by developers as the baseline. Synonyms are generated from this baseline to form the initial population $P$ of size $n$.
  2. 2.
    Generation Operators: Randomly selecting one or two individuals from $P$, new individuals are generated and added to a new population $C$ of size $n$.
  3. 3.
    Evaluation: Each individual in $P$ and $C$ is evaluated, and the num of jailbreaks as a fitness value is assigned to each system message in the population.
  4. 4.
    Selection: Finally, $n$ individuals with the lowest fitness values from populations $P$ and $C$ are selected as the new populations for subsequent iterations.
For clarity, a flowchart illustrating the SMEA framework is presented in Figure 2. In addition, we describe the details of SMEA in Algorithm 1.
Algorithm 1 SMEA algorithm
  Input: the long message LM1, the evaluation set PLe, population size $n$, num of generations $g$
  Output: the final population of robust system messages
  Use gpt-3.5-turbo to expand LM1 to get initial population list $P$.
@@ 5.1 SMEA framework
           $p_{1}$ = randomselect($P$)
           $p$ = rephrase($p_{1}$)
           $C$.append($p$)
        end for
@@
  end if
  if generation == ’crossover’ then
     for $t=1$ to $g$ do
        for $i=1$ to $n$ do
           $p_{1}$ = randomselect($P$)
           $p_{2}$ = randomselect($P$)
           $p$ = crossover($p_{1}$, $p_{2}$)
           $C$.append($p$)
@@
5.2 Generation Operators
Due to the text creation advantages of LLMs, we use LLMs for text creation. We have established two distinct methods for generating base system messages:
  • Crossover: We select two system messages from the original population and input them to LLMs, asking it to combine the two system messages to generate a system message.
  • Rephrase: We require LLMs to understand system message and rewrite them to ensure that the semantics remain unchanged.
Details of the prompts for crossover and rephrase are provided in Appendix A Table 6. According to the methods of generating, we call them SMEA-C and SMEA-R respectively. The SMEA is shown in Algorithm 1.
@@ 6 Experiments
Figure 3: The num of jailbreak in models with final population
@@
In this section, we delve into the experimental methodology in detail, providing insights into the outcomes of both SMEA-R and SMEA-C, and scrutinize the effectiveness of the SMEA framework.
We configured the population size to be $20$ individuals, with $30$ iterations for the evolutionary process. Details of our initialized population are outlined in Appendix A Table 7 and Table 8. Notably, we select $200$ jailbreak prompts from the PL as the evaluation set named PLe for evaluation(assigning fitness to individuals) in the process of evolution. The remaining $100$ jailbreak prompts from the PL are named PLt and used as the test set.
Figure Figure 3 illustrates the performance of the final populations generated by both SMEA-C and SMEA-R against the remaining PL, PS, and PN datasets. It’s important to note that the sizes of PS and PN datasets are both $300$, while the PLt as test set consisting of $100$ jailbreak prompts. An analysis of Figure Figure 3 reveals that while the SMEA-R population exhibits low variance, the SMEA-C population displays high variance. However, the populations derived from SMEA-R demonstrate superior resistance to jailbreak prompts.
Moreover, we identify the top three system messages within each population based on the overall number of jailbreaks in PS, PLt and PN. These system messages exhibit the highest resistance to PS, PLt, and PN prompts, as showcased in Table Table 3. The results indicate that through SMEA, we achieve an impressive evasion rate of 98.9% against jailbreak prompts.
Table 3: The best $3$ individuals in SMEA
@@
SMEA-R 4(1.3%) 2(2.0%) 2(0.7%) 8(1.1%)
0(0.0%) 13(13.0%) 0(0.0%) 13(1.9%)
5(1.7%) 1(1.0%) 9(3.0%) 15(2.1%)
SMEA-C 65(21.7%) 22(22.0%) 38(12.7%) 125(17.8%)
95(31.7%) 6(6.0%) 57(19.0%) 158(22.6%)
115(38.3%) 25(25.0%) 69(23.0%) 209(29.9%)
7 Conclusion and Future work
This paper delves into the influence of system messages to jailbreak prompts in LLMs, shedding light on the transferability characteristics of jailbreak prompts. Furthermore, we propose SMEA, a novel approach aimed at search system messages to thwart jailbreak prompts effectively. Experimental results validate the efficacy of our proposed methodology, showcasing its ability to thwart up to 98.9% of jailbreak prompts. Our investigation not only bolsters the security of large language models but also elevates the barrier against potential jailbreak endeavors.
For future work, we aim to delve deeper into the transferability of jailbreak prompts by employing more comprehensive system messages. Additionally, our focus will encompass the development of dynamic system message configurations within the model to counter various forms of jailbreak prompts. Furthermore, we will aim to combine SMEA-R and SMEA-C methodologies to generate populations characterized by reduced variance and enhanced robustness to jailbreak prompts.
Appendix A Appendix
Table 4: System messages 1
System message content
 short system message
’You are a helpful, respectful and honest assistant.’
 long system message

### v2

@@ 4 Jailbreak experiments with different system messages
    First, we apply PS to Models with Long system messages (ML1) and Models with No system messages (MN1), respectively.
@@
    Next, we apply PL to Models with Short system messages (MS1) and MN1, respectively.
@@
    Finally, we apply PN to MS1 and ML1, respectively.
Table 1: Jailbreak results (num and ASR) in GPT3.5-turbo-0613, LLAMA2 (7b-chat, 13b-chat) with different system messages.
GPT3.5-turbo-0613 LLAMA2-7b-chat LLAMA2-13b-chat
MS1 ML1 MN1 MS1 ML1 MN1 MS1 ML1 MN1
PS 114 (38.0%) 242 (80.6%) 41 (13.7%) 173 (57.7%) 171 (57.0%) 268 (89.3%)
PL 251 (83.6%) 273 (91.0%) 103 (34.3%) 109 (36.3%) 291 (97.0%) 203 (67.7%)
PN 50 (16.7%) 41 (13.6%) 93 (31.0%) 21 (7.0%) 275 (91.7%) 98 (32.7%)
all 301 (50.2%) 155 (25.8%) 515 (85.8%) 196 (32.7%) 62 (10.3%) 282 (47.0%) 566 (94.3%) 269 (44.8%) 471 (78.5%)
Table 2: Jailbreak results (num and ASR) in LLAMA2 (7b, 13b) and VICUNA (7b, 13b) with different system messages.
LLAMA2-7b LLAMA2-13b VICUNA-7b VICUNA-13b
MS1 ML1 MN1 MS1 ML1 MN1 MS1 ML1 MN1 MS1 ML1 MN1
PS 161 (53.7%) 197 (65.7%) 106 (35.3%) 285 (95.0%) 244 (81.3%) 276 (92.0%) 291 (97%) 289 (96.3%)
PL 206 (68.7%) 206 (68.7%) 198 (66.0%) 145 (48.3%) 275 (91.7%) 282 (94.0%) 293 (97.7%) 287 (95.7%)
PN 298 (99.3%) 228 (76.0%) 283 (94.3%) 80 (26.7%) 273 (91.0%) 257 (85.7%) 285 (95.0%) 281 (93.7%)
all 504 (84.0%) 389 (64.8%) 403 (67.2%) 481 (80.2%) 186 (31.0%) 430 (71.7%) 548 (91.3%) 501 (83.5%) 558 (93.0%) 578 (96.3%) 572 (95.3%) 576 (96.0%)
The results, as depicted in Table 1 and Table 2, illustrate the frequency of prompts leading to successful jailbreaks. It is important to note that the models in Table 1 are models with RLHF and models in Table 2 without RLHF. By these tables, we can discover that LLMs equipped with different system messages show large differences in ASR for the same jailbreak prompts. For example, LLAMA2-7b-chat equipped with ML1 has only 7% ASR against PN, while LLAMA2-7b-chat equipped with MN1 has 100% ASR against PN. At this point, we can answer RQ1:
Response to RQ1: Observing the PS, PL and PN rows of Table 1 and Table 2, it becomes evident that LLMs, with the exception of VICUNA (7b, 13b), display markedly distinct ASR in response to identical jailbreak prompts when equipped with varying system messages. This finding underscores the pivotal role of system messages in influencing jailbreak occurrences within LLMs.
Additionally, by examining the last row of Table 1 and Table 2, we observe that LLMs equipped with ML1 generally exhibit a lower overall ASR. However, it is noteworthy that LLAMA2-13b-chat with MS1 displays a 97% ASR when encountering PN, whereas LLAMA2-13b-chat with MN1 shows a 67.7% ASR under the same conditions. Furthermore, LLAMA2-13b with MS1 and MN1 demonstrates an unexpectedly large disparity in ASR when tested against PL. At this point, we can answer RQ2:
Response to RQ2: There is some transferability between jailbreak prompts generated for different system messages. Although in certain cases, LLMs equipped with MS exhibit weaker resistance compared to those equipped with MN, LLMs with the ML1 consistently exhibit the lowest ASR, irrespective of whether the LLMs have undergone RLHF.
5 System Messages Evolutionary Algorithm
Given the significant impact of system messages to jailbreaks, we first test whether just making a little change in system messages could change the ASR of LLMs further to verify the possibility of optimizing system messages against jialbreaks. Based on the obtained conclusion, we then propose the System Messages Evolutionary Algorithm (SMEA), devised to procure optimal system messages that mitigate jailbreaks risks.
5.1 Observation of ASR in Synonymous System Messages
To test whether just making a little change in the system messages could change the ASR of LLMs, we generate synonyms sentence that the length of system messages remain essentially unchanged for the above three system messages and the synonymous sentences are shown in Appendix A.4. These messages are then set as system messages to get models MS2, ML2, and MN2, respectively. Subsequently, we employ PS, PL, and PN to jailbreak MS2, ML2, and MN2, and the results are presented in Table 3. Observing each row of Table 3, we discovered that, with the exception of VICUNA (7b, 13b), merely substituting system messages with synonymous sentences of nearly identical length results in significant alterations to the ASR. At this point, we can answer RQ3:
Response to RQ3: With the exception of VICUNA (7b, 13b), making a little change to the system messages could change the ASR of LLMs. This indicates that we can achieve more robust system messages by changing system messages dimly.
Figure 3: The main framework of SMEA. The details of the workflow that is explained in Section 5.2.
5.2 SMEA Framework
Based on answer to RQ3, we propose System Messages Evolutionary Algorithm (SMEA), devised to procure optimal system messages that mitigate jailbreaks risks. The SMEA framework comprises four primary phases: initialization, generation, evaluation, and selection.
  1. 1.
    Initialization: In this phase, we leverage the system message commonly utilized by developers as the initial seed. Similar sentences are generated from the initial seed to form the initial population $P$ of size $n$.
  2. 2.
    Generation: Randomly selecting one or two individuals from parent population $P$, new individuals are generated and added to a child population $C$ of size $n$.
  3. 3.
    Evaluation: Each individual in $P$ and $C$ is evaluated, and the ASR as a fitness value is assigned to each system message in the population.
@@ 5.2 SMEA Framework
    Selection: $n$ individuals with the lowest fitness values from populations $P$ and $C$ are selected as the new populations for subsequent iterations.
For clarity, a flowchart illustrating the SMEA framework is presented in Figure 3.
5.3 Generation Operators
Due to the text generation capabilities of LLMs, we use LLMs for text generation. We have established three distinct methods for generating system messages:
  • Rephrase: We require LLMs to understand system message and rephrase them to ensure that the semantics remain unchanged.
  • Crossover: We randomly select two system messages from population and input them to LLMs, asking it to combine the two system messages to generate a system message.
  • Mixed: We randomly select two system messages from the population for crossover and rephrase the result obtained from the crossover.
Details of the prompts for crossover and rephrase are provided in Appendix A.5. According to the methods of generating, we call them SMEA-R, SMEA-C and SMEA-X respectively. The details of SMEA is shown in Algorithm 1.
Table 3: The jailbreak results (num and ASR) for models with synonymous sentences of system messages.
@@ 5.3 Generation Operators
GPT3.5-turbo-0613 203(67.7%) 207(69.0%) 212(70.7%)
LLAMA2-7b 200(66.7%) 281(93.7%) 239(79.7%)
LLAMA2-13b 210(70.0%) 298(99.3%) 257(85.7%)
LLAMA2-7b-chat 88(29.3%) 197(65.7%) 138(46.0%)
LLAMA2-13b-chat 228(76.0%) 300(100%) 234(78.0%)
VICUNA-7b 288(96.0%) 290(96.7%) 288(96.0%)
VICUNA-13b 280(93.3%) 292(97.3%) 294(98.0%)
Algorithm 1 System Messages Evolutionary Algorithm
  Input: the long message LM1, the evaluation set PLe, population size $n$, num of generations $g$, method (rephrase, crossover, mixed)
  Output: the final population of robust system messages $P$
  Use GPT3.5-turbo to expand LM1 to get initial population list $P$.
@@
        $p_{1}$ = random_select($P$)
        if method == ’rephrase’ then
           $C$.append(rephrase($p_{1}$))
        else if method == ’crossover’ then
           $p_{2}$ = random_select($P$)
           $C$.append(crossover($p_{1}$, $p_{2}$))
        else if method == ’mixed’ then
           $p_{2}$ = random_select($P$)
           $C$.append(rephrase(crossover($p_{1}$, $p_{2}$)))
        end if
@@ 6 Experiments
In this section, we describe the experimental in detail, providing insights into the outcomes of both SMEA-R, SMEA-C and SMEA-X and scrutinize the effectiveness of the SMEA framework.
Considering the cost of computation, we set the population size to be $20$, with $30$ iterations for the evolutionary process. Details of our initialized population are outlined in Appendix A.5. Notably, we select $200$ jailbreak prompts of PL as the evaluation set named PLe for evaluation in the process of evolution. This is because Table 1 and Table 2 demonstrate that system messages capable of resisting PL can also effectively resist PS and PN. The remaining $100$ jailbreak prompts of PL are named PLt and used as the test set. In addition, we select GPT3.5-turbo as text generation LLM and the temperature parameter is set to be $1$. In order to analyze the performance and stability differences between SMEA-R, SMEA-C, SMEA-X, we illustrates the combined performance of the final populations generated by both SMEA-C, SMEA-R and SMEA-X against the remaining datasets (PLt, PS, and PN) in Appendix A.2.
Observing the result of Appendix A.2, we discover that:
  1. 1.
    SMEA-R always shows the smallest variance as well as the worst performance. This may be because ’Rephrase’ makes each individual in the population does not change much semantically, yet ’Crossover’ and ’Mixed’ are more likely to produce sentences that are semantically different from the original system messages.
  2. 2.
    The performance of SMEA-C and SMEA-X are uneven, which may be because in ‘Mixed’, sometimes the system message generated through ’Crossover’ performs better than the one generated through ’Crossover’ and ’Rephrase’, but this individual is not saved.
  3. 3.
    Consistent with the analysis in Table 3, except for VICUNA (7b, 13b), using SMEA to optimize system messages can increase the resistance of LLMs to jailbreak prompts above 60%. However, for the VICUNA (7b, 13b), the minimum ASR still remains as high as 41.4%. To examine the impact of the SMEA on VICUNA (7b, 13b), we charted the evolutionary trajectory of VICUNA (7b, 13b) in Appendix A.2. Appendix A.2 illustrates that while the ASR declines at a slow rate, there is an overarching downward trend throughout the whole process. There are two reasons for this:
    • The number of iterations and the size of the population in algorithm set too small.
    • VICUNA (7b, 13b) itself is more susceptible to jailbreak than other LLMs.
Based on the above observations, we can answer RQ4:
Response to RQ4: Leveraging the text generation capabilities of LLMs in conjunction with the principles of evolutionary algorithms, presents a significant advancement in crafting system messages that exhibit a high degree of diversity and robustness. This method parallels the process of natural selection, where the most effective system messages are retained and refined over successive generations, thereby enhancing the LLM’s ability to resist and counteract jailbreak prompts.
Finally, in Appendix A.1, we select the top-3 system messages in each population based on performance in PS, PLt and PN.
7 Conclusion
This paper delves into the influence of system messages to jailbreak prompts in LLMs, answering the question of whether the longer system message is more robust. In addition, we answer whether the ASR of LLMs change significantly when there is a little change (similar in semantics, a little change in length) in system messages. Furthermore, we propose SMEA, a novel approach aimed at search system messages to thwart jailbreak prompts effectively. Our investigation not only bolsters the security of large language models but also elevates the barrier against potential jailbreak endeavors.
8 Limitations
The first limitation of SMEA lies in the potential for the population to get trapped in local optima during the optimization process. This issue partly arises from the initialization of the population with system messages. Moreover, our designed prompts may not fully exploit the language generation capabilities of LLMs to produce a more diverse range of natural language outputs. In the future, exploring diverse population initialization strategies and optimizing prompts to enhance performance remains a promising area for further research.
The second limitation of SMEA is related to experimental resources. Due to resource constraints, jailbreak prompts separately for LLMs equipped with different system messages are not enough and we are unable to jailbreak with sufficient amount of system messages. This limitation had an impact on the evaluation during the iterative process. Additionally, compared to evolutionary algorithms, SMEA defaults to only $20$ iterations in our experimental setup. The smaller number of iterations and limited evaluation resources somewhat underestimate the performance potential of SMEA.
Appendix A Appendix
A.1 The Top-3 System messages of LLMs
We select the top-3 system messages in each population based on the ASR in PS, PLt and PN.
Table 4: The top-3 results in SMEA of GPT3.5-turbo-0613
@@ A.1 The Top-3 System messages of LLMs
SMEA-C 4 (1.3%) 2 (2.0%) 2 (0.7%) 8 (1.1%)
0 (0.0%) 13 (13.0%) 0 (0.0%) 13 (1.9%)
5 (1.7%) 1 (1.0%) 9 (3.0%) 15 (2.1%)
SMEA-R 65 (21.7%) 22 (22.0%) 38 (12.7%) 125 (17.8%)
95 (31.7%) 6 (6.0%) 57 (19.0%) 158 (22.6%)
115 (38.3%) 25 (25.0%) 69 (23.0%) 209 (29.9%)
SMEA-M 2 (0.7%) 0 (0.0%) 4 (1.3%) 6 (0.9%)
3 (1.0%) 3 (3.0%) 1 (0.3%) 7 (1.0%)
4 (1.3%) 5 (5.0%) 2 (0.7%) 11 (1.6%)
Table 5: The top-3 results in SMEA of llama2-7b
PS PLt PN all
SMEA-C 0 (0.0%) 0 (0.0%) 1 (0.3%) 1 (0.1%)
0 (0.0%) 1 (1.0%) 1 (0.3%) 2 (0.3%)
0 (0.0%) 1 (1.0%) 16 (5.3%) 17 (2.4%)
SMEA-R 2 (0.7%) 0 (0.0%) 41 (13.7%) 43 (6.1%)
2 (0.7%) 3 (3.0%) 41 (13.7%) 46 (6.6%)
4 (1.3%) 4 (4.0%) 38 (12.7%) 46 (6.6%)
SMEA-M 0 (0.0%) 2 (2.0%) 3 (1.0%) 5 (0.7%)
0 (0.0%) 0 (0.0%) 18 (6.0%) 18 (2.6%)
1 (0.3%) 11 (11.0%) 9 (3.0%) 21 (3.0%)
Table 6: The top-3 results in SMEA of LLAMA2-7b-chat
PS PLt PN all
SMEA-C 1 (0.3%) 0 (0.0%) 2 (0.7%) 3 (0.4%)
8 (2.7%) 1 (1.0%) 5 (1.7%) 14 (2.0%)
14 (4.7%) 2 (2.0%) 9 (3.0%) 25 (3.6%)
SMEA-R 65 (21.7%) 30 (30.0%) 25 (8.3%) 120 (17.1%)
60 (20.0%) 28 (28.0%) 38 (12.7%) 126 (18.0%)
65 (21.7%) 32 (32.0%) 31 (10.3%) 128 (18.3%)
SMEA-M 41 (13.7%) 8 (8.0%) 21 (7.0%) 70 (10.0%)
54 (18.0%) 12 (12.0%) 31 (10.3%) 97 (13.9%)
50 (16.7%) 22 (22.0%) 28 (9.3%) 100 (14.3%)
Table 7: The top-3 results in SMEA of LLAMA2-13b
PS PLt PN all
SMEA-C 0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
SMEA-R 102 (34.0%) 67 (67%) 68 (22.7%) 237 (33.9%)
105 (35.0%) 68 (68.0%) 68 (22.7%) 241 (34.4%)
101 (33.7%) 75 (75.0%) 67 (22.3%) 243 (34.7%)
SMEA-M 0 (0.0%) 0 (0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
Table 8: The top-3 results in SMEA of LLAMA2-13b-chat
PS PLt PN all
SMEA-C 0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
SMEA-R 100 (33.3%) 73 (73.0%) 93 (31.0%) 266 (38.0%)
100 (33.3%) 72 (72.0%) 97 (32.3%) 269 (38.4%)
100 (33.3%) 72 (72.0%) 97 (32.3%) 269 (38.4%)
SMEA-M 0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
Table 9: The top-3 results in SMEA of VICUNA-7b
PS PLt PN all
SMEA-C 192 (64.0%) 57 (57.0%) 210 (70.0%) 459 (65.6%)
188 (62.7%) 65 (65.0%) 225 (75.0%) 478 (68.3%)
192 (64.0%) 59 (59.0%) 235 (78.3%) 486 (69.4%)
SMEA-R 263 (87.7%) 93 (93.0%) 237 (79.0%) 593 (84.7%)
265 (88.3%) 94 (94.0%) 244 (81.3%) 603 (86.1%)
270 (90.0%) 97 (97.0%) 241 (80.3%) 608 (86.9%)
SMEA-M 181 (60.3%) 59 (59.0%) 216 (72.0%) 456 (65.1%)
187 (62.3%) 59 (59.0%) 224 (74.7%) 470 (67.1%)
192 (64.0%) 59 (59.0%) 236 (78.7%) 487 (69.6%)
Table 10: The top-3 results in SMEA of VICUNA-13b
PS PLt PN all
SMEA-C 97 (32.3%) 70 (70.0%) 123 (41.0%) 290 (41.4%)
107 (35.7%) 74 (74.0%) 165 (55.0%) 346 (49.4%)
104 (34.7%) 80 (80.0%) 168 (56.0%) 352 (50.3%)
SMEA-R 106 (35.3%) 74 (74.0%) 161 (53.7%) 341 (48.7%)
139 (46.3%) 84 (84.0%) 173 (57.7%) 396 (56.6%)
161 (53.7%) 79 (79.0%) 182 (60.7%) 422 (60.3%)
SMEA-M 266 (88.7%) 87 (87.0%) 269 (89.7%) 622 (88.9%)
264 (88.0%) 94 (94.0%) 267 (89.0%) 625 (89.3%)
269 (89.7%) 89 (89.0%) 268 (89.3%) 626 (89.4%)
A.2 The Visualization of Final Population
We present the performance metrics for the final population in Figure 4. Additionally, the evolutionary trajectories of VICUNA (7b, 13b) are illustrated in Figure LABEL:vicunatre.
Figure 4: The ASR of LLMs in final populations. In this figure, we present the performance on the final populations obtained from various generative methods across different LLMs.
A.3 The System Messages
In Table 11, we give the system messages used in our experiments.
Table 11: The system messages
System messages Content
short system message ’You are a helpful, respectful and honest assistant.’